OpenStudy (anonymous):

what is the y-intercept of f(x)= 2x+2i

OpenStudy (solomonzelman):

x-y=5 -y=5-x y=-5+x y=x+5 See what I am doing ?

OpenStudy (solomonzelman):

Now, can you tell me the slope?

OpenStudy (anonymous):

1?

OpenStudy (solomonzelman):

Yes, there we go :)

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

you welcome !